# Continuous cloud position spectroscopy using a magneto-optical trap

### v1

@@ Acknowledgements.
We acknowledge helpful discussions with M. Borkowski and E. Baaij. We thank SURF for maintaining the White Rabbit link between VSL and UvA. We also thank R.J.C. Spreeuw, Mikhail Volkov and Florian Schäfer for useful discussions and review of this manuscript. This work has received funding from the European Union’s (EU) Horizon 2020 research and innovation program under Grant Agreement No. 820404 (iqClock project) and No. 860579 (MoSaiQC project). It also received funding from the European Partnership on Metrology, co-financed from the European Union’s Horizon Europe Research and Innovation Programme and by the Participating States, under project 22IEM01 TOCK. It further received funding from the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme.
@@ A.3 Temperature Limit in a Five-Beam MOT
provided that $v_{\text{6,cap}}>v_{\text{5,cap}}$. Experimental results [38] show that capture velocities can easily exceed $v_{\text{5,cap}}$ be an order of magnitude when atoms enter the MOT region counter-propagating the bottom beam (i.e., beam against gravity direction). In this regime, the bottom beam effectively acts as a white-light and Zeeman slower. In steady state, the MOT temperature determines the typical thermal velocity, $v_{\text{rms}}=\sqrt{3k_{B}T/m}$, which must remain well below $v_{\text{5,cap}}$ to ensure confinement. This leads to the condition
@@ A.5 Shift Measurements for Error Budget
Figure 6: Sensitivity of the MOT position to bias magnetic fields. The position shift along the gravity (z) axis is measured while scanning bias coil currents along the x, y, and z directions, with a magnetic field gradient of $0.39~\mathrm{G/cm}$. A linear shift of $29.86(2)~\mu\mathrm{m}/\mathrm{mG}$ is observed for fields along the z-direction (red triangles). For fields along the x (orange stars) and y (light red crosses) directions, the position shift follows a quadratic dependence $z(B)=A\times(B-B_{\text{offset}})^{2}$, with fitted amplitudes A=0.0096(4)A=-0.0096(4) and $0.0095(2)~\mu\mathrm{m}/\mathrm{mG}^{2}$, and offsets $B_{\text{offset}}=391.88(1.59)$ and $117.15(19)~\mathrm{mG}$, respectively.
As shown in Fig. 6, we observe the strongest effect for a magnetic field applied along the vertical direction. Here, we measure a linear shift of $29.86(2)~\mu\mathrm{m}/\mathrm{mG}$ over a 400 mG range (red triangles). In contrast, applying a bias magnetic field along the x- and y-directions leads to a more subtle effect. While one might not expect a change in the MOT’s vertical position from horizontal fields, lateral displacement causes the atoms to experience different intensity regions of the bottom MOT beam due to its Gaussian intensity profile. This alters the local radiation pressure, requiring the MOT to shift vertically to reach a new equilibrium. The resulting position shift is well described by a quadratic dependence of the form $z(B)=A\times(B-B_{\text{offset}})^{2}$. We extract similar amplitudes in both directions: $A=-0.0096(4)~\mu\mathrm{m}/\mathrm{mG}^{2}$ in x, and $0.0095(2)~\mu\mathrm{m}/\mathrm{mG}^{2}$ in y. This agreement is consistent with the symmetry of the Gaussian beam profile in the transverse plane. However, the fitted offsets differ significantly: $B_{\text{offset}}=391.88(1.59)~\mathrm{mG}$ in x, and $117.15(0.19)~\mathrm{mG}$ in y. These differences suggest that the center of the beam profile is not perfectly aligned with the quadrupole magnetic field center. Centering would improve stability by rendering vertical MOT position changes only quadratically dependent on horizontal field changes. Overall, the vertical (z-direction) bias magnetic field induces the largest shift, dominating the magnetic contribution to the systematic error. Shifts induced by bias fields in the transverse directions are more than an order of magnitude smaller, though still measurable and relevant for high-precision applications.
@@
Figure 7: Sensitivity of the MOT position to atom number changes. The atom number is varied by loading the MOT for a finite time before imaging. All measurements are performed at a magnetic field gradient of $0.33~\mathrm{G/cm}$. The main plot shows a linear dependence of the MOT position on atom number with a slope of $29.84(2.71)~\mu\mathrm{m}/(10^{8}~\text{atoms})$. Inset: correlation between the in situ MOT cloud waist and vertical position, used to infer atom number changes during continuous operation. A linear fit yields a slope of 0.76(8)-0.76(8).
@@
Figure 7 shows that increasing atom number results in a downward shift of the MOT’s equilibrium position. We observe a linear sensitivity of $29.84(2.71)~\mu\mathrm{m}/(10^{8}~\text{atoms})$. This behavior is attributed to increased optical density at higher atom numbers, which leads to partial shielding of the atomic cloud from the vertical MOT beam. As a result, the radiation pressure is reduced, and the MOT shifts downward to a new equilibrium position where the local Zeeman shift compensates the force imbalance. Since precise control of atom number is challenging in our system, we developed a method to infer and compensate for atom number fluctuations during continuous operation. Specifically, we use the vertical MOT waist size measured from absorption images as a proxy for the MOT atom number. We confirm that this size also correlates with the MOT position (see inset of Fig. 7), and we assume a similar correlation holds for fluorescence imaging. We propose this method instead of relying on total fluorescence counts, since at the high optical densities found in our BB MOT such counts lead to inaccurate atom number estimates. The measured sensitivity is 0.76(8)-0.76(8). This technique therefore provides a means to correct for atom number fluctuations, relaxing the stability requirement on atom number by approximately an order of magnitude.
@@
The third noise source we examine is intensity fluctuations in the MOT beams. Here, it is important to distinguish between fluctuations in the horizontal beams and in the single vertical beam. To characterize their respective influence on MOT position, we again use blue absorption imaging and measure the MOT’s vertical displacement while varying the power of the two types of beams. All measurements are performed at a magnetic field gradient of $0.33~\mathrm{G/cm}$. Starting with the vertical beam, we vary its power from approximately 80% to 100% of its nominal value and observe a positive linear dependence of the MOT position on beam power. The measured slope is $7.25(55)~\mu\mathrm{m}/\%$. This behavior is consistent with expectations: an increase in vertical beam power leads to a stronger radiation pressure from below. To restore force balance, the MOT shifts upward, requiring a smaller local Zeeman shift to reach equilibrium. In contrast, fluctuations in the horizontal beam power result in a negative position shift, with a slope of $1.41(23)~\mu\mathrm{m}/\%$. This effect arises from increased compression of the MOT at higher horizontal beam intensities, which in turn increases atomic density. The resulting increase in optical density leads to partial shielding of the vertical beam, reducing the effective light force and requiring the MOT to shift downward to reestablish equilibrium. Both vertical and horizontal power fluctuations could, in the future, be stabilized below the 1% level using active power control, making their contribution to long-term position drift manageable.
@@
Figure 8: Sensitivity of the MOT position to power fluctuations in the horizontal and vertical MOT beams. The position shift along the vertical (gravity) direction is measured while changing the power of each type of MOT beam. All measurements are taken at a magnetic field gradient of $0.33~\mathrm{G/cm}$. For the horizontal beams (red triangles), we observe a linear shift of $1.41(23)~\mu\mathrm{m}/\%$, while for the vertical beam (orange stars), the shift is $7.25(55)~\mu\mathrm{m}/\%$.

### v2

@@ Acknowledgements.
We acknowledge helpful discussions with M. Borkowski and E. Baaij. We thank SURF for maintaining the White Rabbit link between VSL and UvA. We also thank R.J.C. Spreeuw, Mikhail Volkov and Florian Schäfer for useful discussions and review of this manuscript. This work has received funding from the European Union’s (EU) Horizon 2020 research and innovation program under Grant Agreement No. 820404 (iqClock project) and No. 860579 (MoSaiQC project). It also received funding from the European Partnership on Metrology, co-financed from the European Union’s Horizon Europe Research and Innovation Programme and by the Participating States, under project 22IEM01 TOCK. It further received funding from the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. The raw data and the analysis tools used in this manuscript can be found in Reference [11].
@@ A.3 Temperature Limit in a Five-Beam MOT
provided that $v_{\text{6,cap}}>v_{\text{5,cap}}$. Experimental results [8] show that capture velocities can easily exceed $v_{\text{5,cap}}$ by an order of magnitude when atoms enter the MOT region counter-propagating the bottom beam (i.e., beam against gravity direction). In this regime, the bottom beam effectively acts as a white-light and Zeeman slower. In steady state, the MOT temperature determines the typical thermal velocity, $v_{\text{rms}}=\sqrt{3k_{B}T/m}$, which must remain well below $v_{\text{5,cap}}$ to ensure confinement. This leads to the condition
@@ A.5 Shift Measurements for Error Budget
Figure 6: Sensitivity of the MOT position to bias magnetic fields. The position shift along the gravity (z) axis is measured while scanning bias coil currents along the x, y, and z directions, with a magnetic field gradient of $0.39~\mathrm{G/cm}$. A linear shift of $29.9(2)~\mu\mathrm{m}/\mathrm{mG}$ is observed for fields along the z-direction (red triangles). For fields along the x (orange stars) and y (light red crosses) directions, the position shift follows a quadratic dependence $z(B)=A\times(B-B_{\text{offset}})^{2}$, with fitted amplitudes A=0.0096(4)A=-0.0096(4) and $0.0095(2)~\mu\mathrm{m}/\mathrm{mG}^{2}$, and offsets $B_{\text{offset}}=391.9(1.6)$ and $117.2(0.2)~\mathrm{mG}$
, respectively.
As shown in Fig. 6, we observe the strongest effect for a magnetic field applied along the vertical direction. Here, we measure a linear shift of $29.9(2)~\mu\mathrm{m}/\mathrm{mG}$ over a 400 mG range (red triangles). In contrast, applying a bias magnetic field along the x- and y-directions leads to a more subtle effect. While one might not expect a change in the MOT’s vertical position from horizontal fields, lateral displacement causes the atoms to experience different intensity regions of the bottom MOT beam due to its Gaussian intensity profile. This alters the local radiation pressure, requiring the MOT to shift vertically to reach a new equilibrium. The resulting position shift is well described by a quadratic dependence of the form $z(B)=A\times(B-B_{\text{offset}})^{2}$. We extract similar amplitudes in both directions: $A=-0.0096(4)~\mu\mathrm{m}/\mathrm{mG}^{2}$ in x, and $0.0095(2)~\mu\mathrm{m}/\mathrm{mG}^{2}$ in y. This agreement is consistent with the symmetry of the Gaussian beam profile in the transverse plane. However, the fitted offsets differ significantly: $B_{\text{offset}}=391.9(1.6)~\mathrm{mG}$ in x, and $117.2(0.2)~\mathrm{mG}$ in y. These differences suggest that the center of the beam profile is not perfectly aligned with the quadrupole magnetic field center. Centering would improve stability by rendering vertical MOT position changes only quadratically dependent on horizontal field changes. Overall, the vertical (z-direction) bias magnetic field induces the largest shift, dominating the magnetic contribution to the systematic error. Shifts induced by bias fields in the transverse directions are more than an order of magnitude smaller, though still measurable and relevant for high-precision applications.
@@
Figure 7: Sensitivity of the MOT position to atom number changes. The atom number is varied by loading the MOT for a finite time before imaging. All measurements are performed at a magnetic field gradient of $0.33~\mathrm{G/cm}$. The main plot shows a linear dependence of the MOT position on atom number with a slope of $29.8(2.7)~\mu\mathrm{m}/(10^{8}~\text{atoms})$. Inset: correlation between the in situ MOT cloud waist and vertical position, used to infer atom number changes during continuous operation. A linear fit yields a slope of 0.76(8)-0.76(8).
@@
Figure 7 shows that increasing atom number results in a downward shift of the MOT’s equilibrium position. We observe a linear sensitivity of $29.8(2.7)~\mu\mathrm{m}/(10^{8}~\text{atoms})$. This behavior is attributed to increased optical density at higher atom numbers, which leads to partial shielding of the atomic cloud from the vertical MOT beam. As a result, the radiation pressure is reduced, and the MOT shifts downward to a new equilibrium position where the local Zeeman shift compensates the force imbalance. Since precise control of atom number is challenging in our system, we developed a method to infer and compensate for atom number fluctuations during continuous operation. Specifically, we use the vertical MOT waist size measured from absorption images as a proxy for the MOT atom number. We confirm that this size also correlates with the MOT position (see inset of Fig. 7), and we assume a similar correlation holds for fluorescence imaging. We propose this method instead of relying on total fluorescence counts, since at the high optical densities found in our BB MOT such counts lead to inaccurate atom number estimates. The measured sensitivity is 0.76(8)-0.76(8). This technique therefore provides a means to correct for atom number fluctuations, relaxing the stability requirement on atom number by approximately an order of magnitude.
@@
The third noise source we examine is intensity fluctuations in the MOT beams. Here, it is important to distinguish between fluctuations in the horizontal beams and in the single vertical beam. To characterize their respective influence on MOT position, we again use blue absorption imaging and measure the MOT’s vertical displacement while varying the power of the two types of beams. All measurements are performed at a magnetic field gradient of $0.33~\mathrm{G/cm}$. Starting with the vertical beam, we vary its power from approximately 80% to 100% of its nominal value and observe a positive linear dependence of the MOT position on beam power. The measured slope is $7.3(6)~\mu\mathrm{m}/\%$. This behavior is consistent with expectations: an increase in vertical beam power leads to a stronger radiation pressure from below. To restore force balance, the MOT shifts upward, requiring a smaller local Zeeman shift to reach equilibrium. In contrast, fluctuations in the horizontal beam power result in a negative position shift, with a slope of $1.4(2)~\mu\mathrm{m}/\%$. This effect arises from increased compression of the MOT at higher horizontal beam intensities, which in turn increases atomic density. The resulting increase in optical density leads to partial shielding of the vertical beam, reducing the effective light force and requiring the MOT to shift downward to reestablish equilibrium. Both vertical and horizontal power fluctuations could, in the future, be stabilized below the 1% level using active power control, making their contribution to long-term position drift manageable.
@@
Figure 8: Sensitivity of the MOT position to power fluctuations in the horizontal and vertical MOT beams. The position shift along the vertical (gravity) direction is measured while changing the power of each type of MOT beam. All measurements are taken at a magnetic field gradient of $0.33~\mathrm{G/cm}$. For the horizontal beams (red triangles), we observe a linear shift of $1.4(2)~\mu\mathrm{m}/\%$, while for the vertical beam (orange stars), the shift is $7.3(6)~\mu\mathrm{m}/\%$.